an object slide down an inclined plane from an initial height of 30m its velocity at the foot of the plane is 20ms, the percentage of its initial potential energy which is clssinated as heat is
1
Expert's answer
2021-10-19T15:13:03-0400

The initial potential energy of an object

"E_p=mgh"

The final kinetic energy of the object

"E_k=\frac{mv^2}{2}"

The amount of mechanical energy transred into heat

"\frac{E_p-E_k}{E_p}100\%=\frac{mgh-mv^2/2}{mgh}100\%\\
=(1-v^2/2gh)100\%\\=(1-20^2/(2*9.8*30))*100\%=32\%"